OpenStudy (anonymous):

please help me Solve -2x^2 + 8x - 8 = 0

OpenStudy (solomonzelman):

1) divide everything in the equation by -2 2) complete the square Hint: x²-4x+4=(x-2)² 3) finish solving Let me know if you need more help.

OpenStudy (anonymous):

OMG I don't know this stuff.

OpenStudy (solomonzelman):

I am going to give a similar example \[3x^2+18x+27=0\]\[3(x^2+6x+9)=0\]\[x^2+6x+9=0\]\[(x+3)^2=0\]\[(x+3)=0\]\[x=-3\]

OpenStudy (p0sitr0n):

-2x^2+8x-8=0 _________________ -2 x^2-4x+4=0 x^2-4x+4=(x-2)^2=0 x-2=0 x=2

OpenStudy (anonymous):

I'm suppose to graduate next Thursday and just found out I don't have a passing grade in Alg II. I'm extremely stressed and have been crying all day. I'm frustrated.... I'm about to try it.